Perimeter of a rectangle. What is the width of a rectangular playground that is
2x-5 feet and the length is 3x+9 feet. Can you please write a polynomial P(x) that represents the perimeter and then evalute this perimeter polynomial if x is 4 feet? Thanks!

Found 2 solutions by checkley77, KADINE:
Answer by checkley77(12569)

p(x)(2x-5)(3x+9)
p(4)=(2*4-5)(3*4+9)
p(4)=(8-5)(12+9)
p(4)=3*21
p(4)=63 ans.

Answer by KADINE(2)

P(*)(28*-5)(3*+9)P(4)=(*4-5)(3*+9)P(4)=(8-5)(12+9)P(4)=3*21)P=36